# BSTcontrol

### v2

@@ 3.1.2 Transceiver Design
Transceiver design represents a pivotal area in wireless communication, and large AI models are increasingly demonstrating their capability to transform traditional methodologies. Conventional transceiver designs often rely on handcrafted algorithms tailored to specific channel models. However, the adaptability and complexity of large AI models offer a paradigm shift, enabling data-driven approaches that can learn and optimize transceiver functionalities directly from communication data. This encompasses various aspects of transceiver design, from channel coding and decoding to modulation and demodulation, and even end-to-end system optimization. Recent surveys, such as the review on machine learning for channel coding in [113], highlight the burgeoning interest in AI-driven channel coding techniques. Furthermore, research into AI-aided receivers, exemplified by [114] on adaptive orthogonal frequency division multiplexing receivers, showcases the practical application and performance gains achievable through AI in receiver design.
@@
A key paradigm enabled by WLAM is the end-to-end optimization of the entire communication system. This approach treats the transmitter and receiver as components of a single neural network, allowing for joint optimization of their parameters using gradient back-propagation. The central challenge in this framework is the physical wireless channel itself. Because the channel introduces stochastic noise and distortions in a way that is not directly differentiable with respect to the transmitter output, it prevents direct gradient flow from the receiver output back to the transmitter input. To enable end-to-end training via back-propagation, a differentiable path between transmitter and receiver is required during the training phase. WLAM offers solutions to this. For instance, a differentiable channel surrogate model can be learned explicitly or implicitly using techniques like diffusion models [95] or generative adversarial networks (GANs). This learned, differentiable model mimics the channel behavior while allowing gradient flow, acting as a bridge during offline training. Alternatively, the system can be trained using a known, mathematically differentiable channel model (such as additive white Gaussian noise or a simplified fading model that simulates noisy conditions), similar to training an autoencoder [115]. This end-to-end optimization allows the system to learn communication strategies potentially superior to traditional designs constrained by modular optimization.
@@
Transformers, originally developed for natural language processing, now exemplify one powerful architecture being adapted within this end to end optimization framework. A novel approach is introduced in [115] using transformers as generative models to reduce channel error rates in end-to-end transceivers . This work demonstrates the capability of transformers to learn complex error correction codes and optimize the entire transceiver chain. Further investigation into transformer architectures for channel decoding is presented in [116] which proposes a linear transformer architecture to efficiently decode 5G low-density parity-check (LDPC) codes, achieving competitive performance with reduced computational complexity. These studies indicate that transformers, with their powerful attention mechanisms, can capture intricate relationships within communication signals and offer a promising avenue for designing advanced AI-native transceivers.
@@ 3.1.3 Integrated Sensing and Communication
The next generation communication system is expected to achieve both high-rate data transmission and high-accuracy sensing. Traditional methods often design communication and sensing systems separately, which is becoming inadequate for meeting these dual requirements simultaneously, especially given the increasing scarcity of spectrum resources [118]. Integrated sensing and communication (ISAC) emerges as a crucial paradigm, aiming to reuse spectrum and hardware for both functionalities. However, conventional ISAC designs often struggle with the complexities of real-world environments and the integration of rich contextual information. Specifically, handling multimodal sensing data, which could significantly enhance both communication and sensing performance, poses a significant challenge for traditional ISAC systems that are typically tailored for unimodal data processing.
Large AI models, particularly multimodal LLMs (MLLMs), offer an effective approach to address these limitations. MLLMs, trained on massive multimodal datasets, possess the ability to deeply understand and integrate semantically complex multimodal information [96]. By leveraging MLLMs, ISAC systems can move beyond unimodal operation and effectively fuse data from diverse sensors, such as RGB-D cameras, LiDAR, and radar, to create a more comprehensive and nuanced understanding of the environment. This enhanced environmental perception capability facilitates sensing-assisted communication by enabling more precise channel modeling, proactive blockage prediction, and adaptive beamforming [96, 119]. Conversely, communication-assisted sensing benefits from MLLMs through improved data sharing and collaborative perception among distributed agents, such as unmanned aerial vehicle (UAV) swarms, leading to enhanced sensing coverage and accuracy. Furthermore, MLLMs can be employed for multi-objective optimization in ISAC systems, balancing the trade-offs between communication and sensing performance, as demonstrated in UAV networks by [119]. The integration of MLLMs into ISAC systems thus promises to unlock a new era of intelligent wireless systems capable of simultaneously delivering high-performance communication and sophisticated environmental perception.
@@ 3.1.4 Wireless Caching
Wireless caching, a technique to store frequently accessed data closer to users, is crucial for reducing latency and improving network efficiency in wireless communications [120, 121]. Traditional wireless caching relies on static algorithms and heuristics for data placement and retrieval. These methods often struggle to adapt to the dynamic nature of wireless traffic, diverse user demands, and the sheer volume of content in modern networks. To overcome these limitations, large AI models are emerging as a promising approach to revolutionize wireless caching strategies [122].
@@
The integration of large AI models into wireless caching systems holds the potential to significantly enhance wireless communication performance. By dynamically adapting caching strategies based on learned patterns and predictions, large AI models can minimize data retrieval latency, reduce backhaul traffic, and improve overall throughput. Moreover, the flexible nature of large AI models allows for optimization across various caching objectives. By adjusting the training objectives, large AI model-driven caching systems can be tailored for energy efficiency [123], delay reduction [124], throughput maximization [125], or a combination of these, thereby creating highly adaptable and performant wireless caching solutions for next-generation networks.
@@ 3.1.5 Large AI Model for Physical Layer Security
Large AI models are rapidly transforming wireless communications, but this advancement necessitates a critical reassessment of physical layer security (PLS). While AI promises enhanced performance, the inherent complexity of these models, especially generative AI, introduces new security vulnerabilities that must be addressed for robust 6G networks [126]. Traditional PLS techniques, while valuable, often lack the adaptability and sophistication to counter threats in AI integrated wireless systems. Generative AI, on the other side, offers a powerful paradigm shift, providing tools to both analyze and enhance security at the physical layer.
A comprehensive security framework for WLAM based PLS must incorporate robust threat modeling [127]. This involves identifying potential attack vectors targeting the AI models themselves within the physical layer context. For example, adversarial attacks could manipulate wireless signals to deceive WLAM based receivers, classifiers, or channel estimators. Attackers might also attempt model poisoning during distributed training phases or aim to infer sensitive information by analyzing model outputs or parameters. Countermeasures leveraging WLAM capabilities include adversarial training to enhance model resilience against perturbed inputs, utilizing generative models like GANs or diffusion models to detect anomalies deviating from learned normal channel or signal behavior, and developing robust aggregation methods in federated learning scenarios to mitigate poisoning threats [126]. Establishing such a technical framework is crucial for designing and evaluating secure WLAM deployments.
@@
Furthermore, the dynamic nature of wireless environments and threats necessitates adaptive security mechanisms. RL presents a promising approach for enabling WLAM to dynamically adjust PLS strategies in response to real time conditions. An RL agent could monitor the communication environment, observing network state information, estimated channel quality, and outputs from threat detection modules. Based on this state, the RL agent could decide actions such as adjusting the secret key generation rate in secret key generation protocols, selecting optimal wiretap coding schemes, modifying artificial noise injection levels, or adapting parameters of authentication protocols like physical unclonable functions challenge response mechanisms [128]. The reward function for the RL agent could balance security metrics, such as achieved secrecy rates or successful authentication counts, against communication performance indicators like throughput and latency. This allows the system to intelligently trade off security and efficiency based on the current context and detected threat level.
@@
The MoE architecture offers an efficient structure for implementing such adaptive security [97]. MoE-enabled generative AI frameworks can enhance standalone models by reducing computational complexity and improving adaptability. By combining multiple specialized expert models that each trained for a specific task (e.g., jamming resilience, eavesdropping mitigation), MoE offers flexible, context-aware security solutions. A gating mechanism, informed by an RL agent or environmental sensing, dynamically selects the most appropriate expert based on the detected scenario. This architecture facilitates targeted defenses, enhances scalability, and improves adaptability to evolving wireless security challenges, integrating well with the dynamic adjustments driven by RL. The ongoing research combining MoE, generative AI, and RL promises a new generation of intelligent and adaptive PLS for future wireless networks.
@@ 3.2.1 Network Operations and Management
LLMs provide powerful capabilities for enhancing wireless network operations, including real time monitoring, sophisticated fault diagnosis, and proactive security protection [98, 129]. Leveraging their advanced natural language processing abilities, LLMs can analyze diverse and voluminous data sources common in wireless environments, such as base station logs, radio access network (RAN) performance metrics, user mobility data, and security alerts. They can identify subtle anomalies indicative of wireless specific issues like interference, handover failures, or configuration drifts in RAN components. By learning patterns from historical data, LLMs can predict potential faults, enabling preemptive actions to maintain network stability and service quality [98]. Furthermore, they can assist network administrators by diagnosing problems through interactive dialogues, interpreting complex technical information, and generating concise reports or suggesting resolution steps, thereby improving operational efficiency, particularly in complex multi vendor or Open RAN (O-RAN) settings.
@@ 3.2.2 Network Optimization and Resource Allocation
Optimizing performance in wireless networks involves allocating scarce resources like spectrum, power, and computational capabilities across dynamic channel conditions and user mobility. LLMs offer intelligent solutions for these network layer optimization tasks [98]. They can analyze network state information and user requirements to make informed decisions regarding radio resource allocation, dynamic spectrum allocation, network slice configuration for diverse QoS demands (e.g., URLLC, eMBB, mMTC in 6G), and traffic routing in complex topologies involving terrestrial and non terrestrial segments [98, 130]. Frameworks utilizing LLMs, potentially adapted using efficient techniques like data driven low rank adaptation, demonstrate potential in optimizing tasks such as adaptive bitrate streaming or job scheduling in edge computing environments, outperforming conventional algorithms by better understanding complex system dynamics and user behavior [99]. This facilitates automated, context aware optimization that adapts to the fluid nature of wireless environments.
@@ 3.2.3 Network Configuration Automation
LLMs can automate and enhance the design and configuration of wireless networks [129]. They can assist engineers by translating high level service requirements or natural language intents into specific network configurations, such as RAN parameter settings, cell planning parameters, or network slice definitions [131]. LLMs can generate configuration scripts or code for various network functions and protocols, potentially reducing manual effort and minimizing errors. Their ability to reason over complex dependencies, possibly augmented by external knowledge bases or verification tools, allows them to validate configurations, detect potential conflicts (e.g., policy violations, incorrect parameter settings), and ensure consistency across network elements, which is particularly valuable in increasingly complex and disaggregated architectures like O-RAN [129, 131]. LLMs can also be used to generate test scripts for validating wireless software systems, using synthetically generated test data that mirrors real world network conditions [132].
@@ 3.2.4 AI Native Architectures for Wireless Networks
The integration of LLMs drives the evolution towards AI native network architectures, where intelligence is embedded throughout the network rather than being an overlay [130]. These architectures envision LLMs operating collaboratively across cloud, edge, and potentially device layers. Frameworks like NetGPT [100] exemplify this by using smaller, specialized LLMs at the wireless edge for tasks like prompt enhancement and context personalization based on local information (e.g., user location, device state), while leveraging larger cloud based LLMs for complex generative tasks, enabling efficient personalized services . Other frameworks such as NetLLM focuses on creating adaptable LLM based systems capable of handling diverse networking tasks through structured workflows involving components for analysis, planning, calculation, and interaction with network tools and environments [99, 131, 129]. These architectures are fundamental to realizing the 6G vision of intelligent, automated, and highly flexible communication systems.
@@ 3.2.5 Adaptation and Integration Techniques for Wireless LLMs
Bridging the gap between general purpose LLMs and the specific demands of the wireless network layer requires sophisticated adaptation and integration methods. Handling the unique data modalities in wireless communication, such as time series signal data, CSI, graph based topology information, or specific protocol formats, necessitates multimodal encoders capable of projecting this diverse information into a space understandable by the LLM [99, 129]. PEFT techniques, like LoRA, are critical for instilling domain specific knowledge into LLMs without the prohibitive cost of full retraining, making adaptation feasible even for resource constrained edge nodes [99, 100]. Advanced prompt engineering, including CoT and RAG drawing upon telecom standards or operational manuals, guides LLMs towards accurate reasoning and factual responses [131, 129]. Crucially, seamless integration with external network tools, simulators, controllers, and verifiers is essential to enable LLMs to not just reason about the network, but also actively participate in its management and control.

### v4

@@ 3.1.2 Transceiver Design
Transceiver design represents a pivotal area in wireless communication, and large AI models are increasingly demonstrating their capability to transform traditional methodologies. Conventional transceiver designs often rely on handcrafted algorithms tailored to specific channel models. However, the adaptability and complexity of large AI models offer a paradigm shift, enabling data-driven approaches that can learn and optimize transceiver functionalities directly from communication data. This encompasses various aspects of transceiver design, from channel coding and decoding to modulation and demodulation, and even end-to-end system optimization. Recent surveys, such as the review on machine learning for channel coding in [accessdecoding], highlight the burgeoning interest in AI-driven channel coding techniques. Furthermore, research into AI-aided receivers, exemplified by [TWCDecode] on adaptive orthogonal frequency division multiplexing receivers, showcases the practical application and performance gains achievable through AI in receiver design.
@@
A key paradigm enabled by WLAM is the end-to-end optimization of the entire communication system. This approach treats the transmitter and receiver as components of a single neural network, allowing for joint optimization of their parameters using gradient back-propagation. The central challenge in this framework is the physical wireless channel itself. Because the channel introduces stochastic noise and distortions in a way that is not directly differentiable with respect to the transmitter output, it prevents direct gradient flow from the receiver output back to the transmitter input. To enable end-to-end training via back-propagation, a differentiable path between transmitter and receiver is required during the training phase. WLAM offers solutions to this. For instance, a differentiable channel surrogate model can be learned explicitly or implicitly using techniques like diffusion models [10104549] or generative adversarial networks (GANs). This learned, differentiable model mimics the channel behavior while allowing gradient flow, acting as a bridge during offline training. Alternatively, the system can be trained using a known, mathematically differentiable channel model (such as additive white Gaussian noise or a simplified fading model that simulates noisy conditions), similar to training an autoencoder [choukroun2022error]. This end-to-end optimization allows the system to learn communication strategies potentially superior to traditional designs constrained by modular optimization.
@@
Transformers, originally developed for natural language processing, now exemplify one powerful architecture being adapted within this end to end optimization framework. A novel approach is introduced in [choukroun2022error] using transformers as generative models to reduce channel error rates in end-to-end transceivers. This work demonstrates the capability of transformers to learn complex error correction codes and optimize the entire transceiver chain. Further investigation into transformer architectures for channel decoding is presented in [hernandez20255g] which proposes a linear transformer architecture to efficiently decode 5G low-density parity-check (LDPC) codes, achieving competitive performance with reduced computational complexity. These studies indicate that transformers, with their powerful attention mechanisms, can capture intricate relationships within communication signals and offer a promising avenue for designing advanced AI-native transceivers.
@@ 3.1.3 Integrated Sensing and Communication
The next generation communication system is expected to achieve both high-rate data transmission and high-accuracy sensing. Traditional methods often design communication and sensing systems separately, which is becoming inadequate for meeting these dual requirements simultaneously, especially given the increasing scarcity of spectrum resources [liu2022integrated]. Integrated sensing and communication (ISAC) emerges as a crucial paradigm, aiming to reuse spectrum and hardware for both functionalities. However, conventional ISAC designs often struggle with the complexities of real-world environments and the integration of rich contextual information. Specifically, handling multimodal sensing data, which could significantly enhance both communication and sensing performance, poses a significant challenge for traditional ISAC systems that are typically tailored for unimodal data processing.
Large AI models, particularly multimodal LLMs (MLLMs), offer an effective approach to address these limitations. MLLMs, trained on massive multimodal datasets, possess the ability to deeply understand and integrate semantically complex multimodal information [cheng2025large]. By leveraging MLLMs, ISAC systems can move beyond unimodal operation and effectively fuse data from diverse sensors, such as RGB-D cameras, LiDAR, and radar, to create a more comprehensive and nuanced understanding of the environment. This enhanced environmental perception capability facilitates sensing-assisted communication by enabling more precise channel modeling, proactive blockage prediction, and adaptive beamforming [cheng2025large, li2025large]. Conversely, communication-assisted sensing benefits from MLLMs through improved data sharing and collaborative perception among distributed agents, such as UAV swarms, leading to enhanced sensing coverage and accuracy. Furthermore, MLLMs can be employed for multi-objective optimization in ISAC systems, balancing the trade-offs between communication and sensing performance, as demonstrated in UAV networks by [li2025large]. The integration of MLLMs into ISAC systems thus promises to unlock a new era of intelligent wireless systems capable of simultaneously delivering high-performance communication and sophisticated environmental perception.
@@ 3.1.4 Wireless Caching
Wireless caching, a technique to store frequently accessed data closer to users, is crucial for reducing latency and improving network efficiency in wireless communications [paschos2016wireless, 10292740, 10606215]. Traditional wireless caching relies on static algorithms and heuristics for data placement and retrieval. These methods often struggle to adapt to the dynamic nature of wireless traffic, diverse user demands, and the sheer volume of content in modern networks. To overcome these limitations, large AI models are emerging as a promising approach to revolutionize wireless caching strategies [sheraz2020artificial].
@@
The integration of large AI models into wireless caching systems holds the potential to significantly enhance wireless communication performance. By dynamically adapting caching strategies based on learned patterns and predictions, large AI models can minimize data retrieval latency, reduce backhaul traffic, and improve overall throughput. Moreover, the flexible nature of large AI models allows for optimization across various caching objectives. By adjusting the training objectives, large AI model-driven caching systems can be tailored for energy efficiency [jiang2019deep], delay reduction [tsai2018mobile], throughput maximization [cheng2018localized], or a combination of these, thereby creating highly adaptable and performant wireless caching solutions for next-generation networks.
@@ 3.1.5 Large AI Model for Physical Layer Security
Large AI models are rapidly transforming wireless communications, but this advancement necessitates a critical reassessment of physical layer security (PLS). While AI promises enhanced performance, the inherent complexity of these models, especially generative AI, introduces new security vulnerabilities that must be addressed for robust 6G networks [tccnsecurity]. Traditional PLS techniques, while valuable, often lack the adaptability and sophistication to counter threats in AI integrated wireless systems. On the other side, Generative AI offers a powerful paradigm shift, providing tools to both analyze and enhance security at the physical layer.
A comprehensive security framework for WLAM based PLS must incorporate robust threat modeling [10110346]. This process involves identifying and formalizing potential attack vectors that target the AI models themselves within the physical layer context. Examples of such threats include adversarial attacks, where manipulated wireless signals deceive WLAM-based receivers, and model poisoning, where malicious data is injected during distributed training phases to corrupt the model.
To counter these threats, the capabilities of large AI models themselves can be leveraged to create powerful defense mechanisms. These countermeasures include adversarial training to enhance model resilience against perturbed inputs, the use of generative models to detect anomalies deviating from normal signal behavior, and the development of robust aggregation methods in federated learning to mitigate poisoning threats [tccnsecurity]. Establishing a technical framework that integrates both threat models and their corresponding countermeasures is crucial for designing and evaluating secure WLAM deployments.
@@
Besides, the dynamic nature of wireless environments and threats necessitates adaptive security mechanisms. RL presents a promising approach for enabling WLAM to dynamically adjust PLS strategies in response to real time conditions. An RL agent could monitor the communication environment, observing network state information, estimated channel quality, and outputs from threat detection modules. Based on this state, the RL agent could decide actions such as adjusting the secret key generation rate in secret key generation protocols, selecting optimal wiretap coding schemes, modifying artificial noise injection levels, or adapting parameters of authentication protocols like physical unclonable functions challenge response mechanisms [mitev2023physical]. The reward function for the RL agent could balance security metrics, such as achieved secrecy rates or successful authentication counts, against communication performance indicators like throughput and latency. This allows the system to intelligently trade off security and efficiency based on the current context and detected threat level.
@@
Furthermore, the MoE architecture offers an efficient structure for implementing such adaptive security [wcmsecurity]. MoE-enabled generative AI frameworks can enhance standalone models by reducing computational complexity and improving adaptability. By combining multiple specialized expert models that are each trained for a specific task (e.g., jamming resilience, eavesdropping mitigation), MoE offers flexible, context-aware security solutions. A gating mechanism, informed by an RL agent or environmental sensing, dynamically selects the most appropriate expert based on the detected scenario. This architecture facilitates targeted defenses, enhances scalability, and improves adaptability to evolving wireless security challenges, integrating well with the dynamic adjustments driven by RL. The ongoing research combining MoE, generative AI, and RL promises a new generation of intelligent and adaptive PLS for future wireless networks.
@@ 3.2.1 Network Operations and Management
LLMs provide powerful capabilities for enhancing wireless network operations, including real time monitoring, sophisticated fault diagnosis, and proactive security protection [10829820, liuLLMNetwork]. Leveraging their advanced natural language processing abilities, LLMs can analyze diverse and voluminous data sources common in wireless environments, such as base station logs, radio access network (RAN) performance metrics, user mobility data, and security alerts. They can identify subtle anomalies indicative of wireless specific issues like interference, handover failures, or configuration drifts in RAN components. By learning patterns from historical data, LLMs can predict potential faults, enabling preemptive actions to maintain network stability and service quality [10829820]. Furthermore, they can assist network administrators by diagnosing problems through interactive dialogues, interpreting complex technical information, and generating concise reports or suggesting resolution steps, thereby improving operational efficiency, particularly in complex multi vendor or Open RAN (O-RAN) settings.
@@ 3.2.2 Network Optimization and Resource Allocation
Optimizing performance in wireless networks involves allocating scarce resources like spectrum, power, and computational capabilities across dynamic channel conditions and user mobility. LLMs offer intelligent solutions for these network layer optimization tasks [10829820]. They can analyze network state information and user requirements to make informed decisions regarding radio resource allocation, dynamic spectrum allocation, network slice configuration for diverse QoS demands (e.g., URLLC, eMBB, mMTC in 6G), and traffic routing in complex topologies involving terrestrial and non terrestrial segments [9537935]. Frameworks utilizing LLMs, potentially adapted using efficient techniques like data-driven low-rank adaptation, demonstrate potential in optimizing tasks such as adaptive bitrate streaming or job scheduling in mobile edge computing environments, outperforming conventional algorithms by better understanding complex system dynamics and user behavior [NetLLM]. This facilitates automated, context-aware optimization that adapts to the fluid nature of wireless environments.
Beyond LLMs, generative diffusion models have emerged as powerful tools for exploring complex solution spaces in non-convex optimization. Authors in [11018297] and [11006143] proposed DiffSG and GDSG respectively, leveraging diffusion models as generative solvers to effectively address optimization problems in network management and mobile edge computing. For multi-objective scenarios, [11085045] demonstrated the potential of diffusion models in enhancing monitoring efficiency within UAV-enabled IoT systems. Furthermore, [10839314] systematically explored and analyzed the feasibility of deploying diffusion models as core network optimizers.
@@ 3.2.3 Network Configuration Automation
LLMs can automate and enhance the design and configuration of wireless networks [liuLLMNetwork]. They can assist engineers by translating high level service requirements or natural language intents into specific network configurations, such as RAN parameter settings, cell planning parameters, or network slice definitions [huang2023large]. LLMs can generate configuration scripts or code for various network functions and protocols, potentially reducing manual effort and minimizing errors. Their ability to reason over complex dependencies, possibly augmented by external knowledge bases or verification tools, allows them to validate configurations, detect potential conflicts (e.g., policy violations, incorrect parameter settings), and ensure consistency across network elements, which is particularly valuable in increasingly complex and disaggregated architectures like O-RAN [liuLLMNetwork, huang2023large]. LLMs can also be used to generate test scripts for validating wireless software systems, using synthetically generated test data that mirrors real world network conditions [10615269].
@@ 3.2.4 AI Native Architectures for Wireless Networks
The integration of LLMs drives the evolution towards AI native network architectures, where intelligence is embedded throughout the network rather than being an overlay [9537935]. These architectures envision LLMs operating collaboratively across cloud, edge, and potentially device layers. Frameworks like NetGPT [chen2024netgpt] exemplify this by using smaller, specialized LLMs at the wireless edge for tasks like prompt enhancement and context personalization based on local information (e.g., user location, device state), while leveraging larger cloud based LLMs for complex generative tasks, enabling efficient personalized services. Other frameworks such as NetLLM focuses on creating adaptable LLM based systems capable of handling diverse networking tasks through structured workflows involving components for analysis, planning, calculation, and interaction with network tools and environments [NetLLM, huang2023large, liuLLMNetwork]. These architectures are fundamental to realizing the 6G vision of intelligent, automated, and highly flexible communication systems.
@@ 3.2.5 Adaptation and Integration Techniques for Wireless LLMs
Bridging the gap between general purpose LLMs and the specific demands of the wireless network layer requires sophisticated adaptation and integration methods. Handling the unique data modalities in wireless communication, such as time series signal data, CSI, graph based topology information, or specific protocol formats, necessitates multimodal encoders capable of projecting this diverse information into a space understandable by the LLM [NetLLM, liuLLMNetwork]. PEFT techniques, like LoRA, are critical for instilling domain specific knowledge into LLMs without the prohibitive cost of full retraining, making adaptation feasible even for resource constrained edge nodes [NetLLM, chen2024netgpt]. Advanced prompt engineering, including CoT and RAG drawing upon telecom standards or operational manuals, guides LLMs towards accurate reasoning and factual responses [huang2023large, liuLLMNetwork]. Crucially, seamless integration with external network tools, simulators, controllers, and verifiers is essential to enable LLMs to not just reason about the network, but also actively participate in its management and control.